OpenStudy (anonymous):

How do I factor this? h(x) = x^2 +8x+16. Please help. I came up with this as a perfect square trinomial, but I do not know how to factor it.

OpenStudy (mathstudent55):

A perfect square trinomial can have theses two forms: \(a^2 + 2ab + b^2 = (a + b)^2 \) or \(a^2 - 2ab + b^2 = (a - b)^2 \)

OpenStudy (anonymous):

wolfram alpha it :DD

OpenStudy (mathstudent55):

Since the signs of the x^2 term and the constant term are both positive, as it is in the two forms above, and the sign for the x-term is negative, you have a case of the second form.

OpenStudy (mathstudent55):